OpenStudy (anonymous):

An event lasts 5 hours and 30 minutes. How long did the event last in seconds?

OpenStudy (anonymous):

1 hour = 3600s 1 min = 60 s

OpenStudy (anonymous):

so 30 min = 60 *30 = 1800 s

OpenStudy (anonymous):

19800

OpenStudy (anonymous):

5 hour = 5* 3600 = 18000 so Total = 19800s

OpenStudy (anonymous):

19800 seconds